# From References to Insights: Collaborative Knowledge Minigraph Agents for Automating Scholarly Literature Review

### v1

@@ Introduction
For a detailed comparison, we sample an instance from the Multi-Xscience dataset and use well-known LLMs, GPT-3.5-turbo and GPT 4.0, to generate summaries with the default requirements listed in Table 1. The generated results are shown in Table 4. We find that GPT-3.5-turbo suffers from information loss, omitting citation 13. GPT 4.0 shows improvement but lists facts in parallel without logical connections. For example, citations 37 and 14 are listed side by side, but show no parallel logical relationship. It can even lead to hallucination problems, as no evidence shows citation 16 is a statistical method, while citation 16 and citations 3, 6, 5 are all categorized as statistical methods. We then use different versions of the proposed CKMAs to generate a summary. It can be observed that without the knowledge minigraph construction agent (KMCA), the multiple path summarization agent (MPSA) contributes to highlighting different categories of algorithms from the desired viewpoint. Without MPSA, the KMCA contributes to organizing algorithms logically, e.g., from probabilistic to statistical approaches, and then to example-based learning methods. With both modules, CKMAs generates the best summary, categorizing algorithms as supervised learning as in the gold summary and detailing subcategories in a logical order.
We then analyze the differences between queried public knowledge graphs and the constructed knowledge minigraphs. We sample an instance from the Multi-Xscience dataset for this comparison. To query the knowledge graph, we use SPARQL to access Wikidata, a collaborative knowledge base. The queried knowledge graph is shown in the upper part of Fig. 4. For the knowledge minigraph, we employ the proposed method with knowledge minigraph construction, with the result displayed in the lower part. We observe that the entities in the queried knowledge graph are general-purpose and lack specific insights into research problems. The minigraph clearly presents tasks and methods (some including metrics and materials), making it more informative for summarization purposes.
@@
This paper aims to provide an intelligent research copilot to assist in writing literature reviews based on given references. While recent LLMs excel at natural language understanding and generation, they struggle to explicitly model complex relationships between ideas from multiple papers. To address this challenge, we propose collaborative knowledge minigraph agents (CKMAs). The contributions of this work are threefold: (1) We propose scientific document-oriented knowledge minigraphs and, for the first time, equip LLMs with knowledge minigraphs for multiple scientific document summarization. (2) We are the first to develop a prompt-based iterative algorithm to process a vast amount of literature and automatically construct knowledge minigraphs for multiple scientific document summarization. (3) We firstly introduce a mixture of experts’ mechanisms to attempt the organization of literature reviews with different logical paths on minigraphs and derive the best one via self-evaluation.
We conduct comparison experiments, ablation studies, and case studies. The results show the effectiveness of the proposed method. For future work, we plan to explore fine-tuning LLMs with the proposed CKMAs to better follow instructions and approximate human-written literature reviews in collected datasets. We also intend to investigate the possibility of generating full survey papers with multiple paragraphs, which involve more scientific documents and more complex relationships and require planning of the survey paper’s organization.

### v2

@@ Introduction
For further insights, we sample an instance from the Multi-Xscience dataset and use well-known LLMs, GPT-3.5-turbo and GPT 4.0, to generate summaries. The generated results are shown in Table 4. We find that GPT-3.5-turbo suffers from information loss, omitting citation 13. GPT 4.0 shows improvement but lists facts in parallel without rational connections. For example, citations 37 and 14 are listed side by side, but show no parallel relationship. We then use different versions of the proposed CKMAs to generate a summary. It can be observed that without the KMCA, the MPSA contributes to highlighting different categories of algorithms from the desired viewpoint. Without MPSA, the KMCA contributes to organizing algorithms rationally, e.g., from probabilistic to statistical approaches and then to example-based learning methods. With both modules, CKMAs generate the best summary.
@@
We then analyze the differences between queried public knowledge graphs and the constructed knowledge minigraphs. We sample an instance from the Multi-Xscience dataset for this comparison. To query the knowledge graph, we use SPARQL to access Wikidata, a collaborative knowledge base. The queried knowledge graph is shown in the upper part of Fig. 4. For the knowledge minigraph, we employ the proposed method with knowledge minigraph construction, with the result displayed in the lower part. We observe that the entities in the queried knowledge graph are general-purpose and lack specific insights into research problems. The minigraph clearly presents tasks and methods. It demonstrates GPT-3.5-turbo’s reliability in knowledge minigraph construction, as supported by related studies (Zhu et al. 2024). As a result, knowledge minigraphs provide more topic-focused information that enhances summarization.
Figure 5: Performance comparison of CKMAs with its backbone model (GPT-3.5-turbo) along faithfulness, completeness, and conciseness dimensions. A higher score indicates better performance.
Disscussion
To gain further insights, besides similarity-based metrics (ROUGE), we incorporate FineSurE (Song et al. 2024), a recent LLM-based automated evaluation framework that assesses summarization quality across three dimensions: faithfulness (minimizing factual errors), completeness (covering key facts), and conciseness (avoiding redundant details).
We evaluated the proposed CKMAs against its backbone model on the Multi-Xscience dataset. Fig. 5 demonstrates that CKMAs outperform GPT-3.5-turbo in conciseness and faithfulness, indicating that knowledge minigraphs effectively filter irrelevant relations while retaining viewpoints supported by related work. Both approaches exhibit moderate completeness scores, which can be attributed to the divergent purposes in individual review paragraphs (e.g., discussing feasibility versus discussing limitations).
@@
This paper aims to provide an intelligent research copilot to assist in writing literature reviews based on given references. While recent LLMs excel at natural language understanding and generation, they struggle to explicitly model complex relations in multiple reference documents. To address this challenge, we propose collaborative knowledge minigraph agents (CKMAs).
We conduct comparison experiments, ablation studies, and case studies on benchmark datasets. Experimental results show the effectiveness of the CKMAs, further revealing promising applications of LLMs in scientific research.
Acknowledgements
This work was supported by the “A Large Language Model-powered Clinical Communication Training System Under Cantonese-English Code-switching Scenario” (ITS/397/23), the Hong Kong Research Grants Council (RGC) Theme-based Research Scheme (T43-513/23-N, T41-603/20-R), and the National Natural Science Foundation of China (62472291), with research partially conducted at the Research Institute for Artificial Intelligence of Things (RIAIoT) at PolyU.